Space Stations

NASA ISS On-Orbit Status 1 April 2014

By Marc Boucher
Status Report
April 3, 2014
Filed under , ,
NASA ISS On-Orbit Status 1 April 2014
NASA ISS On-Orbit Status 1 April 2014

Today: Urine Processing Assembly (UPA) Advance Recycle Filter Tank Assembly (ARFTA) Change Out and Shock: Commander (CDR) Wakata and Flight Enginer (FE)-3 Swanson removed ARFTA serial number (S/N) 1 and installed S/N 2.
Swanson then started a series of activities to shock ARFTA S/N 2 using pretreated water from the Waste and Hygiene Compartment (WHC). The shock treatment targeted microbial contamination which needed to be addressed before the ARFTA could be used with the UPA. Following the shock, Swanson swapped the ARFTA tanks to re-install S/N 1.

Burning and Suppression of Solids (BASS)-II Repair: FE-5 Mastracchio completed the BASS-II hardware repair started last Friday. Mastracchio conducted a fan calibration and then placed kapton tape around the outer edge of the BASS-II door, completely sealing the window. Mastracchio completed a fan calibration and conducted airflow readings. Ground teams will analyze the airflow readings against the ones taken prior to the repair last Friday before determining when BASS-II operations can resume. On February 20th, the BASS-II door experienced heat damage when burning a double-sided, 3 mm thick sample in a different flow configuration. The BASS-II investigation examines the burning and extinction characteristics of a wide variety of fuel samples in microgravity. The BASS-II experiment will guide strategies for materials flammability screening for use in spacecraft as well as provide valuable data on solid fuel burning behavior in microgravity. BASS-II results contribute to the combustion computational models used in the design of fire detection and suppression systems in microgravity and on Earth.

Remote Power Controller Module (RPCM) Trips: There were three RPCM trips overnight. They do not appear to be related. LA2B-G RPC 1 and LAS62B-A RPC 7 initially appear to be hybrid Field Event Transistor (FET) events. Initial troubleshooting on these RPCMs was completed. LAD62B-A appears to be an overcurrent event. Teams are meeting to determine follow-on troubleshooting steps for all three RPCM trips.

LA2B-G RPC 1 controls the Lab Moderate Temperature Loop (MTL) regenerative 3-way mix valve. Troubleshooting was performed to close the RPC but it immediately tripped open again. This was repeated 5 times with the same results. To control Lab MTL temperatures, Mastracchio manually moved the regen 3-way mix valve to bypass, then the ground pre-emptively commanded the Lab ITCS to dual mode.
LAS62B-A RPC 7 controls the Lab MTL Nitrogen Introduction Assembly (NIA). Troubleshooting was performed to close the RPC, but it immediately tripped open again. This was repeated 5 times with the same results.
LAD62B-A RPC 12 controls the Carbon Dioxide Removal Assembly (CDRA) valves for the Lab CDRA. The Lab CDRA was in Standby at the time of the trip. No troubleshooting was performed on this RPC today. The next planned operations for the Lab CDRA are this weekend.
Orbital-2 Cygnus Pre-Gather: The crew continued gathering hardware and foam that will be loaded into Cygnus for disposal. These items are being pre-staged in the Permanent Multipurpose Module (PMM) Endcone. Orbital-2 is scheduled to arrive in May.

Today’s Planned Tasks

Reaction Self Test [Complete]
HRF Saliva Ops [Complete]
HRF MELFI Insert [Complete]
SM ??? (Caution & Warning Panel) Test [Complete]
KHROMATOMASS. Experiment Setup / r/g 5220 [Complete]
RSS1,2 Reboot / r/g 1009 [Complete]
KHROMATOMASS. Saliva and Blood Collection / r/g 5220 [Complete]
Laptop RS1(2) Reboot [Complete]
??? Laptops Antivirus Scan Check and Status Report / r/g 5201 [Complete]
Daily Planning Conference (S-band) [Complete]
Recycle tank Changeout [Complete]
IFM UPA Shock Part 1 [Complete]
TOCA WRM Analysis [Complete]
Connecting SM Kurs-P Cables to DC1 port -Y / r/g 1956 [Complete]
KHROMATOMASS SPEKTR M. Photography of the Experiment Ops / r/g 5220 [Complete]
KHROMATOMASS. Closeout Ops / r/g 5220 [Complete]
SHD Daily Questionaire [Complete]
IFM UPA Shock Part 2 [Complete]
Demating ??-7? telemetry connectors. r/g 5202 [Complete]
Replacing ??-7? sensors in Signal-VM system. r/g 5202 [Complete]
IFM NOD2OS5 LHA R&R [Complete]
IFM NOD3FD2 LHA R&R [Complete]
Journal Entry 5
IFM UPA Shock Part 3 [Complete]
ISS adaptation and orientation [Complete]
IFM UPA Shock Part 4 [Complete]
CIR Alignment Guide Install [Complete]
WRM Condensate Pump Init [In work]
54P Stowage Ops with IMS support IMS / r/g 5106 [Complete]
Mating ??-7? sensor telemetry connectors. r/g 5202 [Complete]
IFM UPA Shock Part 5 [Complete]
WHC Piping Retrieval [Complete]
Replacing ??-7? sensors in Signal-VM system – Close-out ops. r/g 5202 [Complete]
??? Maintenance [Complete]
JAXA CM Setup [Complete]
JAXA CM Video [Complete]
IFM UPA Shock Part 6 [Complete]
Photography of Batavia lettuce sample / r/g 5178 [Complete]
JAXA CM Closeout[Complete]
WRM Condensate Pump Term [Complete]
???? Configuration for MRM1 [Complete]
KAPLYA-2. Active Drop Sampler [???] Functionality Test. r/g 5221 [Complete]
IFM UPA Shock Part 7 [Complete]
KAPLYA-2. Active Drop Sampler [???] Functionality Test. r/g 5221 [Complete]
ISS adaptation and orientation [Complete]
IFM UPA Shock Part 8 [Complete]
PAO HD Setup [Complete]
MSG Facility Activation [Complete]
PAO HD Lab Event [Complete]
Node 1 Encoder T/S [Complete]
ISS Crew Handover [Complete]
MRM1 Comm Reconfig [Complete]
24-hour ECG (termination) / r/g 5197
Recycle tank Changeout [Complete]
ISS adaptation and orientation [Complete]
24-hour ECG Recording (start) / r/g 5197 [Complete]
IMS Update [Complete]
BASS Repair Part 1 [Complete]
Cygnus Pregather [Complete]
Deactivation of ?? 712 Gas Analyzer [Complete]
Deactivation of ?? 711 Gas Analyzer. [Complete]
LHA JPM1OF7 R&R [Complete]
EHS TOCA Data Record [Complete]
EHS TOCA WWB Changeout [Complete]
Intermodular TORU Test with Mated ??? 422 (DC1). r/g 5218 [Complete]
BASS Repair Part 2 [Complete]
MSG Facility Power Down [Complete]
VZAIMODEISTVIYE Experiment Run By FE-4 / r/g 4115 [Complete]
Task List

ARED Cylinder Evac
Dragon ROBOT Self Study [In work]
WHC KTO Replace
Roscosmos report
TK-712 GoPro Remove
TK-712 Transfer
Econ photo
SM Printer Swap [Complete]
ARED cycle flywheel reconfig
Cygnus pregather
Food Warmer Setup
HMS Med Kit Photos
IFM LHA JPM1OF1 R&R [Complete]
NORS Prep A/L Reconfig [In work]

Nominal Commanding
Three-Day Look Ahead:

Wednesday, 04/02: ISS Crew Handover & Orientation, EMU Water Sampling, RAM Install
Thursday, 04/03: Hybrid Muscle Training, CFE ICF7, PAO, CIR Bottle Replace
Friday, 04/04: Petri-light measure, ACE Samp, Hybrid Muscle Training,CDRA R&R Prep, Food Warmer Setup
QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – Off
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Off
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.